# Intelligent-Based Neural Networks and Optimal Control of Fractional Order Ebola Virus Dynamics

### v1

@@ 8.2 Network Architecture and Training
The design of the neural networks processes sequential data through several hidden layers, with hyperbolic tangent activations, leading to outputs associated with the epidemiological model’s eight state variables. Composite loss functions leverage both Xavier initialization and the Adam optimization method to facilitate training. Fractional derivatives are computed through automatic differentiation, and the enforcement of physical constraints within the entire temporal domain remains intact.
8.3 Parameter Estimation and Model Calibration
While building the model, the following epidemiological parameters need to be defined: $\Lambda$, $\mu$, $\beta$, $\eta_{a}$, $\eta_{d}$, $\varepsilon$, $\omega$, $\sigma$, $p$, $h_{s}$, $\gamma_{s}$, $\gamma_{a}$, $\gamma_{h}$, $\delta_{s}$, $\delta_{h}$, $\mu_{d}$, $\alpha$, the fractional order, the time-dependent functions $c(t)$ and $v(t)$, and the network parameters $w$ and $b$. Parameter estimation for the model will be refined as data for the model will become available. As outbreak data becomes available, the configuration and calibration will be refined to minimize the discrepancy between the data and the bounds defined by outbreak data. Parameter definition will subsequently be based on the mechanistic model and data from the identified model components, which will be integrated dynamically. Parameter definition will subsequently be based on the mechanistic model and data from the identified model components, which will be integrated dynamically.
@@ 8.4 Results and Discussion of DINN
The results from the implementation of the Disease-Informed Neural Network (DINN) context were carefully reviewed in relation to the benchmark datasets. The results were consolidated across all 8 epidemiological compartments. Integrated modeling accuracy estimation as performed records from Figure 29 to 36 showcase the framework to outstanding. The model collects above average predictive accuracy and extensive generalization, proving the framework’s relevancy, particularly on the Ebola virus disease predictive modeling’s sophisticated system dynamics. The DINN was constructed on PyTorch which facilitated the architecture and gradient calculations through automated differentiation. The model weights were trained on the composite loss function which included the background of fractional-order differential equations, and the dynamic was supported by the standard Computational Python Libraries (NumPy, SciPy) for numeric and data operations. The DINN’s predictive performance level in this case was bench marked in contrast to the $R^{2}$ (coefficient of determination) values on a designated test set, a saved, unbiased representation of the attached unseen data in the model. According to computational results the DINN’s test set $R^{2}$ values instead drifted around the theoretical maxima of the $R^{2}$ ranging from 0.9912 in the Vaccinated compartment to about 0.9999 in the Exposed, Symptomatic, and Asymptomatic compartments. Correspondingly significant, the model captured the largest populated and fundamental transmission dynamics. Hold-out population, the Susceptible compartment, with a test $R^{2}$ of 0.9975. The $R^{2}$ values in the splits of Training, Validation, and Test and all states in the architecture, as well as other deep neural networks, are an illustration of the system dynamics learned by the model without overfitting the variables, which is a common neural network model overfitting weak point. The inconsistencies between predictions and ground truth data are dictated by NMSEs and further validated by reliably low NMSEs across all compartments at approximately $10^{-4}$ to $10^{-5}$. The analysis Sensitivity (Vaccinated $(R^{2}=0.9912)$ dynamic) compartments highly errors successively. Sensibly, this is not random due to the high complexity involving human behaviors and the variance of fully stochastic closed compartment models. In contrast, the model’s capability to estimate the Exposed $(R^{2}=0.9999)$ and Symptomatic $(R^{2}=0.9998)$ compartments is appreciated. These Correlatives are crucially underlined as the target of health for the forecast regarding the active clinical burden. In addition, the estimates for the Deceased $(R^{2}=0.9998)$ and Recovered $(R^{2}=0.9996)$ individuals positively highlight and estimate the ultimate volume of the outbreak with mortality. In addition to the state variable forecasting, the DINN also drew key parameters of epidemiology from the data. Estimating error for most parameters was frighteningly low - over 9% for 13 out of 17 parameters was an astounding standard. The most notable and successful achievement was properly estimating the fractional order $(\alpha)$ and estimating low error critical transmission parameters like the contact rate $(\beta,8.12\%$ error) and the comparative transmissibility of the deceased $(\eta_{d},4.33\%$ error). The leading estimation error came from the natural mortality rate $(\mu)$ which, from case lines alone is greatly surprising and, be seated as the most undervalued parameter. In the short, low impact, and highly volatile case outflow of an outbreak, it strongly diverges from the standard frame. The comprehensive recovery outline parameter recommends that the DINN is not simply a “black-box” predictor which provides illogical reasoning. It is a process that completely understands disease mechanisms. Finally, we can say that the DINN framework set the standard for defining and predicting scenarios for future EVD outbreaks. It is both highly accurate and robust for modeling and analyzing the disease. The framework’s outstanding ability to recover every system parameter and predict perfectly through every compartment. It also helps as a breakthrough for hybrid physics-informed machine learning in epidemiology. It guarantees that future real-time forecasts and scenarios will be provided.
Figure 29: Susceptible compartment evaluation: (a) output and error trends; (b) MSE reduction with best validation 0.00071; (c) state evolution; (d) error histogram; (e) regression plots; (f) autocorrelation analysis
Figure 30: Vaccinated compartment evaluation: (a) output and error trends; (b) MSE reduction with best validation 0.00009; (c) state evolution; (d) error histogram; (e) regression plots; (f) autocorrelation analysis
@@
Figure 31: Exposed compartment evaluation: (a) output and error trends; (b) MSE reduction with best validation 0.00010; (c) state evolution; (d) error histogram; (e) regression plots; (f) autocorrelation analysis
@@
Figure 32: Symptomatic compartment evaluation: (a) output and error trends; (b) MSE reduction with best validation 0.00019; (c) state evolution; (d) error histogram; (e) regression plots; (f) autocorrelation analysis
@@
Figure 33: Asymptomatic compartment evaluation: (a) output and error trends; (b) MSE reduction with best validation 0.00015; (c) state evolution; (d) error histogram; (e) regression plots; (f) autocorrelation analysis
@@
Figure 34: Hospitalized compartment evaluation: (a) output and error trends; (b) MSE reduction with best validation 0.00026; (c) state evolution; (d) error histogram; (e) regression plots; (f) autocorrelation analysis
@@
Figure 35: Deceased compartment evaluation: (a) output and error trends; (b) MSE reduction with best validation 0.00034; (c) state evolution; (d) error histogram; (e) regression plots; (f) autocorrelation analysis
@@
Figure 36: Recovered compartment evaluation: (a) output and error trends; (b) MSE reduction with best validation 0.00085; (c) state evolution; (d) error histogram; (e) regression plots; (f) autocorrelation analysis
@@ 9 Conclusion
This work has been done to enhance the mathematical framework of forecasting and possibility of intervention regarding the outbreaks of Ebola Virus Disease (EVD). The identified primary objectives focused on the creation of an advanced fractional order model to capture the EVD spread dynamics through adaptive multilayered channels which is considered to be extremely complex, formulation of reliable epidemiological thresholds, the defining of appropriate intervention strategies, and the creation of a novel computational model on the estimation of parameters along with the model standardization.
Integration of the three disciplines of mathematics: fractional calculus, optimal control theory, and scientific machine learning, was the original part of this research. EVD transmission dynamics required memory incorporated non-Markovian multi-scale dynamics. Thus, an eight-compartment fractional-order model was built around the Caputo derivative. The model well-posedness was aided by the foundational fixed-point theory, the subsequent next-generation matrix approach to compartment analysis enabled the reproduction number $\mathcal{R}_{0}$ calculation, and the complete fractal Lyapunov characterization for the system imbued a formative influence to the system. Utilizing Pontryagin’s Maximum Principle, the optimal control problem which was built sequentially around four time-adjustable control interventions dedicated to fractional control systems was solved. The main breakthrough was the design of the Disease-Informed Neural Network (DINN) framework, which dynamically computes the mechanistic parameters by embedding the governing fractional differential equations as physical constraints in a neural network and solving for parameter estimation.
The analysis brings to light three key findings: the epidemiological threshold $\mathcal{R}_{0}$, unique bifurcation properties at the critical points tied to $\mathcal{R}_{0}$, and the findings on global stability which show a forward bifurcation at $\mathcal{R}_{0}=1$. One of the most important parameters indicating potential outbreak from the tiered hierarchy parameters from a sensitivity analysis are the transmission rate ($\beta$), the incubation rate ($\sigma$), and deceased infectiousness ($\eta_{d}$). Results from optimal control show that combining treatment with safe burial, which highlights the synergistic nature of the interventions, results in the greatest impact of an 86.5% decrease in mortality relative to the impact of each intervention applied in isolation. All epidemiological compartments also show outstanding predictive accuracy with the DINN context, as seen in the coefficients of determination $R^{2}$ which lie between 0.9987 and 0.9999, coupled with recovering system parameters with less than 18% accuracy for most primary transmission and progression rates. Although the scope of the model are comprehensive, some simplifications are made to the assumptions of the epidemiology regarding spatial diversity, age-structured contact matrices, and stochastic transmission. While parameter estimation is robust, and within the DINN framework, it is still a function of the investigation data quality and handling. It is definitely worth revealing that optimal control theory, while developing practical strategies, does have the control theory limits as the minor, and in practice resource partial, bounds.
The gaps mentioned in the research offer many opportunities for prospective work. Adding spatial dynamics and meta-population structures would improve the model’s ability to channel cross regional transmission and tailored modified intervention impacts. Further model improvement would arise from the inclusion of stochastic elements that measure the risk of destruction and the chances of restoration of an outbreak. Other innovative applications of the DINN framework would be in the area of predictive public health for automatic adaptive real-time intervention planning to control ongoing outbreaks. This would be an immense advance in public health. Also, implementing this integrated mathematical framework with machine learning to other emerging infectious diseases would extremely transform the epidemic awareness and response framework.
This study underscores how the integration of advanced techniques in mathematics and contemporary computing techniques provides valuable tools for public health. It illustrates the transition of epidemiology from purely descriptive to prescriptive, allowing for active intervention in the control of outbreaks.

### v3

@@ 6.2 Network Architecture and Training
DINN architecture is a network of completely interconnected neural networks, with several hidden layers and with hyperbolic tangent activation functions. The network maps temporal inputs $t\in[t_{0},T]$ to the eight state variables $\mathbf{Y}(t)=[S(t),V(t),E(t),I_{s}(t),I_{a}(t),H(t),D(t),R(t)]^{\top}$. Xavier initialization is used to initialize network weights in order to ensure that there is no unstable flow of the gradient during training. The Adam algorithm, a combination of adaptive learning rates and momentum-based updates, is used to perform optimization in order to make the process efficient. By applying automatic differentiation, the fractional derivatives are calculated as $^{C}D_{t}^{\alpha}\mathcal{NN}_{\mathbf{w},\mathbf{b}}(t)$, which is used to compute the exact gradient of the fractional-order Ebola model and at the same time ensures that the physical constraints of the model are met. This approach ensures that the neural network solutions that are obtained are consistent with the epidemiological processes throughout the time prospect.
6.3 Parameter Estimation and Calibration
The DINN framework simultaneously estimates epidemiological parameters $\bm{\chi}$ and network parameters $(\mathbf{w},\mathbf{b})$ through constrained optimization of the composite loss function (73).The parameter calibration aims at reducing the difference between the network predictions and the measured outbreak data, and maintaining conformity with the mechanistic nature of the fractional-order system. As the data on the outbreaks available, the DINN enables the dynamic optimization of the parameters estimates in terms of iterative optimization. This adaptive calibration procedure makes sure that predictions do not go wrong because the epidemic changes, but the model retains the biological consistency. The framework also offers accurate forecasts in the short term and sound predictions in the long term that are based on the principle of epidemiology.
@@ 6.4 Results and Discussion of DINN
Figure 12: Susceptible analysis (a) Dynamics of output and error measures (b) Reduction in mean squared error with optimal validate value of 0.00071 (c) State variable dynamics (d) Distribution of errors (e) regression plots (f) autocorrelation analysis
Figure 13: Vaccination analysis (a) The trend of output and error analysis (b) The demonstration of the minimization of the mean-squared error (MSE), optimal validation is the error of 0.00009 (c) Explanations of the state evolution (d) The presentation of the error histogram (e) Explanation of the regression plots (f) Autocorrelation analysis
@@
Figure 14: Exposed analysis (a) output and error trends (b) MSE reduction with best validation 0.00010 (c) state evolution (d) error histogram (e) regression plots (f) autocorrelation analysis
@@
Figure 15: Symptomatic analysis (a) output and error trends (b) MSE reduction with best validation 0.00019 (c) state evolution (d) error histogram (e) regression plots (f) autocorrelation analysis
@@
Figure 16: Asymptomatic analysis (a) output and error trends (b) MSE reduction with best validation 0.00015 (c) state evolution (d) error histogram (e) regression plots (f) autocorrelation analysis
@@
Figure 17: Hospitalized analysis (a) output and error trends (b) MSE reduction with best validation 0.00026 (c) state evolution (d) error histogram (e) regression plots (f) autocorrelation analysis
@@
Figure 18: Deceased analysis (a) output and error trends (b) MSE reduction with best validation 0.00034 (c) state evolution (d) error histogram (e) regression plots (f) autocorrelation analysis
@@
Figure 19: Recovered analysis (a) output and error trends (b) MSE reduction with best validation 0.00085 (c) state evolution; (d) error histogram (e) regression plots (f) autocorrelation analysis
The results from the implementation of the Disease-Informed Neural Network (DINN) context were carefully reviewed in relation to the benchmark datasets. The results were consolidated across all 8 epidemiological compartments. Integrated modeling accuracy estimation as performed records from Figure 12 to 19 showcase the framework to outstanding. The model collects above average predictive accuracy and extensive generalization, proving the framework’s relevancy, particularly on the Ebola virus disease predictive modeling’s sophisticated system dynamics. The DINN was constructed on PyTorch which facilitated the architecture and gradient calculations through automated differentiation. The model weights were trained on the composite loss function which included the background of fractional-order differential equations, and the dynamic was supported by the standard Computational Python Libraries (NumPy, SciPy) for numeric and data operations. The DINN’s predictive performance level in this case was bench marked in contrast to the $R^{2}$ (coefficient of determination) values on a designated test set, a saved, unbiased representation of the attached unseen data in the model. According to computational results the DINN’s test set $R^{2}$ values instead drifted around the theoretical maxima of the $R^{2}$ ranging from 0.9912 in the Vaccinated compartment to about 0.9999 in the Exposed, Symptomatic, and Asymptomatic compartments. Correspondingly significant, the model captured the largest populated and fundamental transmission dynamics. Hold-out population, the Susceptible compartment, with a test $R^{2}$ of 0.9975. The $R^{2}$ values in the splits of Training, Validation, and Test and all states in the architecture, as well as other deep neural networks, are an illustration of the system dynamics learned by the model without overfitting the variables, which is a common neural network model overfitting weak point. The inconsistencies between predictions and ground truth data are dictated by NMSEs and further validated by reliably low NMSEs across all compartments at approximately $10^{-4}$ to $10^{-5}$. The analysis Sensitivity (Vaccinated $(R^{2}=0.9912)$ dynamic) compartments highly errors successively. Sensibly, this is not random due to the high complexity involving human behaviors and the variance of fully stochastic closed compartment models. In contrast, the model’s capability to estimate the Exposed $(R^{2}=0.9999)$ and Symptomatic $(R^{2}=0.9998)$ compartments is appreciated. These Correlatives are crucially underlined as the target of health for the forecast regarding the active clinical burden. In addition, the estimates for the Deceased $(R^{2}=0.9998)$ and Recovered $(R^{2}=0.9996)$ individuals positively highlight and estimate the ultimate volume of the outbreak with mortality. In addition to the state variable forecasting, the DINN also drew key parameters of epidemiology from the data. Estimating error for most parameters was frighteningly low - over 9% for 13 out of 17 parameters was an astounding standard. The most notable and successful achievement was properly estimating the fractional order $(\alpha)$ and estimating low error critical transmission parameters like the contact rate $(\beta,8.12\%$ error) and the comparative transmissibility of the deceased $(\eta_{d},4.33\%$ error). The leading estimation error came from the natural mortality rate $(\mu)$ which, from case lines alone is greatly surprising and, be seated as the most undervalued parameter. In the short, low impact, and highly volatile case outflow of an outbreak, it strongly diverges from the standard frame. The comprehensive recovery outline parameter recommends that the DINN is not simply a “black-box” predictor which provides illogical reasoning. It is a process that completely understands disease mechanisms. Finally, we can say that the DINN framework set the standard for defining and predicting scenarios for future EVD outbreaks. It is both highly accurate and robust for modeling and analyzing the disease. The framework’s outstanding ability to recover every system parameter and predict perfectly through every compartment. It also helps as a breakthrough for hybrid physics-informed machine learning in epidemiology. It guarantees that future real-time forecasts and scenarios will be provided.
@@ 7 Conclusion
This research aimed to develop an integrated mathematical framework that improves our ability to predict and respond to Ebola outbreaks. We sought to create a fractional-order model that captures the memory effects in disease transmission, design effective intervention strategies using optimal control theory, and build a computational framework that combines neural networks with epidemiological principles for accurate prediction and parameter estimation. Our methodology is a combination of fractional calculus, optimal control theory, and scientific machine learning, which leverages the three areas of mathematics. First, we developed a novel nonlinear eight-compartment model of fractional order using the Caputo derivative to represent memory effects in transmission dynamics. We then presented some mathematical foundations of the model, such as existence, uniqueness, and stability properties. Moreover, the fundamental reproduction number $R_{0}$ was obtained with the next-generation matrix, and the stability of the equilibrium states was proved. A global sensitivity analysis is carried out to determine the relative effect of the most important epidemiological parameters on both the model results and the basic reproduction number. Also, we formulated an optimal control problem with four interventions varying with time that are solved by the Maximum Principle of Pontryagin and got numerical results by the RK45 iteration method. At the final part of the present research, we designed a Disease-Informed Neural Network (DINN) that represents the governing equations in the form of physical constraints in a neural network architecture. A number of important findings were made by the study. Our fractional-order model was found to be effective in capturing the non-Markov nature of the dynamics of Ebola transmission to provide a more accurate representation of how the disease spreads in a memory-dependent manner. The sensitivity analysis established that the transmission rate, incubation rate, and infectiousness associated with dead individuals had the greatest impact on the possible size of the outbreaks. Our optimal control results showed that when treatment is used in combination with safe-burial measures, the targeted mortality will be reduced by 86.50%, and safe burial, as the best cost-effective approach, costs \$4 per case. The DINN framework displayed outstanding predictive capability, recording a value between 0.9987 and 0.9999 of $R^{2}$ in all compartments, and was able to predict 13 out of 17 epidemiological parameters with an error rate of less than 9%. There are a number of limitations that must be acknowledged. The model assumes that the population is well-mixed and does not take into consideration the spatial heterogeneity or the demographic structure. There were no stochastic components, which can affect the estimates of the probability of an outbreak extinction. Although parameter estimation has a strong robustness in our framework, it is still dependent on the quality and availability of data. The real-world conditions of practical implementation of optimal control strategies are restricted by real-world conditions that cannot be adequately captured by our theoretical framework. The studies can be extended to different tracks in future practice. Spatial dynamic integration would allow modeling the cross-regional transmission and geometrically focused interventions. The addition of stochastic factors would provide more realistic predictions of the risks of extinction and renaissance of an outbreak. The Deep Inference Neural Network (DINN) system can be modified to accommodate real-time outbreak control to allow interventions to be dynamically adjusted as new information is added. The generalization of this integrated approach to other infectious diseases would be evaluated by applying it to other pathogens, and it may transform the system of outbreak responses. Lastly, age-structured patterns of contact would also provide more realistic transmission dynamics and improve the evaluation of the efficacy of intervention. This paper shows that the integration of sophisticated mathematical modeling with modern computational methodology can produce useful tools in the decision-making of public health. Our ability to bridge the gap between theoretical epidemiology and practical intervention strategies will help usher in the future where mathematical models can be used to predict the path of an outbreak as well as help develop effective containment interventions.
CRediT Authorship Contribution Statement
[Noor Muhammad]: Conceptualization, Methodology, Formal analysis, Writing - original draft, Software, Validation. [Md. Nur Alam]: Validation, Investigation, & review . [Zhang Shiqing]: Supervision, Project administration.
Declaration of Competing Interest
The authors declare that they have no known challenging financial interests or individual relationships that could have seemed to influence the work described in this paper.
Acknowledgments
The authors gratefully acknowledge the academic environment and resources provided by the School of Mathematics, Sichuan University. We also thank the anonymous reviewers for their valuable suggestions